40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Message ID: 660     Entry time: Mon Apr 15 13:22:53 2019
Author: Alena 
Type: General 
Category: Measurements 
Subject: S1600693 S1600694 S1600695 S1600696 

2019-04-13

  • Laser was clipping on a laminated paper on the optical table. this caused signal on QPD2 and QPD4 be almost 0. Restarted the measurement after the issue was resolved
    • S1600693 in CR1
    • S1600694 in CR2
    • S1600695 in CR3
    • S1600696 in CR4
ELOG V3.1.3-